OpenStudy (anonymous):

[(8 × 4) − 10]^2 + 9

OpenStudy (anonymous):

if your simplifying then its 493

OpenStudy (anonymous):

laura do you know how to solve it??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know the BODMAS rule

OpenStudy (anonymous):

B- Bracket O- Over D- Divide M- Multiply A- Add S- Subtract

OpenStudy (unklerhaukus):

\[[(8 × 4) − 10]^2 + 9\]\[=[32 − 10]^2 + 9\]\[=[22]^2+9\]\[=484+9\]\[=493\]

OpenStudy (anonymous):

also O can be replaced by E- Exponential

OpenStudy (anonymous):

ok now i get it

OpenStudy (anonymous):

in your problem you are having \[\large{[(8\times4)-10]^2+9}\]

OpenStudy (anonymous):

so first do the bracket operations so first bracket operation would be \[\large{[(8\times4)-10]^2 + 9}\] \[=\large{[(32)-10]^2 + 9}\] Now remove the bracket so you will have \[=\large{[32-10]^2 + 9}\] now do the large brackets operation \[=\large{[22]^2 + 9}\] \[=\large{484 + 9}\] \[=\large{493}\]

OpenStudy (anonymous):

ok now i get it